OpenStudy (anonymous):

Find the slope of the line passing through (1, 5) and (0, 2)

OpenStudy (callisto):

Slope \[=\frac{y_2-y_1}{x_2-x_1} = \frac{5-2}{1-0} =?\]

OpenStudy (anonymous):

I dont know i would appreciate it if you could just help me instead of asking me questions that i dont get!

OpenStudy (callisto):

I'm sorry.. Basically, slope is change of y coordinates divided by change of x coordinates. So, we can rewrite it as \[slope = \frac{y_2-y_1}{x_2-x_1}\] Where (x1, y1) is the coordinates of a point and (x2, y2) is another pair of coordinates of the other point Since you're given two points, just plug in the coordinates into the formula above. Do you understand so far?

OpenStudy (anonymous):

yeah kind of

OpenStudy (callisto):

From my first comment here, I've put (x2, y2) = (1, 5) and (x1, y1) =(0, 2) into the formula. Understand it?

OpenStudy (anonymous):

Ok so whats the answer and half and half

OpenStudy (callisto):

It's what I've put a question mark there. All you need to do is to simplify the fraction. Can you try?

OpenStudy (anonymous):

I dont know how okay im stupid so please give me a break on this one please

OpenStudy (callisto):

First, do subtraction on both denominator and numerator, what do you get?

OpenStudy (anonymous):

I dont know its really late and i just need to know real fast and i promise ill try to work it out after

OpenStudy (callisto):

For denominator, it's 1-0 =? For numerator, it's 5-2 =?

OpenStudy (callisto):

Don't worry.. it's fast to solve :)

OpenStudy (anonymous):

ok, 1-0=0 and 5-2=3

OpenStudy (callisto):

Good, now, you get \[slope = \frac{3}{1}\] Can you further simplify it?

OpenStudy (anonymous):

No sorry i dont know how

OpenStudy (callisto):

Wait... 1-0 is not 0 :S

OpenStudy (callisto):

1-0 = 1 :|

OpenStudy (anonymous):

ohh ahaha then 1

OpenStudy (callisto):

anything divided by 1 is its original value. For instance , 2/1 =2 Now, it's 3/1, which is equal to ...?

OpenStudy (anonymous):

3

OpenStudy (callisto):

Yes. There it is! It's the slope you need :)

OpenStudy (anonymous):

Oh so the final answer is 3?

OpenStudy (callisto):

I think it is!